Sign in
Please select an account to continue using cracku.in
↓ →
India plays two matches each with West Indies and Australia.In any match the probabilities of India getting points 0, 1 and 2 are 0.45, 0.05 and 0.50 respectively. Assuming that the outcomes are independent, the probability of India getting at least 7 points is
Total matches played = 4
Maximum points = 2 * 4 = 8
So to get atleast 7 points , India needs to win atleast 3 matches and one can be draw.
1. India wins all 4 matches: There is only one possible outcome
Hence, the probability of 4 wins = (0.5)*(0.5)*(0.5)*(0.5) = 0.0625
2. India wins 3 matches and draws 1: probability in that case = (0.5)*(0.5)*(0.5)*(0.05) = 0.00625
(since 4 cases, any one match an be draw)
P = 0.0625 + 0.025 = 0.0875
Was this solution helpful?
Create a FREE account and get:
Book Free CAT Mentorship
Get personalized CAT strategy from a 99%iler
500+ students mentored
OTP Verification
Enter the 6-digit code sent to your phone
Booking Summary
Enter OTP
Didn't receive the OTP?
Start your IIM journey with the right preparation and crack CAT 2026.
Educational materials for CAT preparation